A ball is drawn at random from a box containing 8 red balls, 2 white balls and 9 blue balls

Determine the probability that the ball drawn is

  1. Red
  2. White
  3. Blue
  4. Not red
  5. Red or white
Answer #1

a) Red

Red = 8; Total balls = 8+2+9 = 19

So the probability of Red is,

8/19 = 0.4211

b) White

2/19 = 0.1053

C) Blue

9/19 = 0.4737

d) Not Red

probability of Red is 0.4211

So Not Red probability is =1 - 0.4211 = 0.5789

e) Red or White

REd = 8 and White = 2

so 10/19 = 0.5263